• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Rikki

    Log entry time 16:08:07 on July30,2003

    Entry number 108208

    keyword=day shift target log

    Begin shift with 140 psig 3He reference cell

    reference cell measurements:
    08:18 vent 3He
    08:31 fill reference cell with 140 psig N2
    08:43 vent/fill reference cell to 90 psig N2
    08:53 vent reference cell to 35 psig N2
    09:21 vent N2 and evacuate cell

    NMR & EPR measurements:
    09:21 move to pick-up coils
    NMR = 41.13%
    EPR = 40.13%

    production running - kin 5.10:
    09:44 move to 3He target

    10:15 adaqlr3 locked up again (same happened yesterday) - rebooted

    continue production running - kin 5.10:
    10:21 turned off lasers 4, 6, 7
    rotated target field to 180 degrees (oops - started rotating to 270 first, then realized JP said 180. After rotation to 270 was complete, rotated back to 0, then to 180 - no choice for rotation from 270 to 180!)
    moved target half-wave plate IN
    10:40 turned on laser 4
    10:41 turned on laser 6
    10:42 turned on laser 7

    production running - kin 5.11:
    11:09 turned off lasers 4, 6, 7
    rotate to 0 degrees
    11:11 moved target half-wave plate OUT
    11:17 turned on laser 4
    11:18 turned on laser 6
    11:19 turned on laser 7
    13:08 turned off lasers 4, 6, 7
    rotate to 270 degrees
    13:12 turn on laser 1
    13:12 turn on laser 2
    13:13 turn on laser 3

    beam studies - rotate target to 0 degrees
    15:08 turn off lasers 1, 2, 3
    rotate to 0 degrees
    15:12 turn on laser 4
    15:13 turn on laser 6
    15:14 turn on laser 7

    15:25 NMR measurement = 41.28%
    15:40 EPR measurement = 39.1%